# Cherry picking in forests: A new characterization for the unrooted hybrid number of two phylogenetic trees

### v3

@@ 6. Bounding the weight of a cherry picking sequence from above by the hybrid number
Claim 7: We can and will assume that neither ${\mathcal{F}}$ nor ${\mathcal{F}}^{\prime}$ contains a cherry in the form of an isolated edge $\{x,y\}$ with $x,y\in L({\mathcal{B}})$.
@@
Proof of Claim 7: Suppose that ${\mathcal{T}}$ is a cherry in ${\mathcal{F}}$ in the form of the isolated edge $\{x,y\}$ with $x,y\in L({\mathcal{B}})$. Let $u,v,\gamma$ and ${\mathcal{N}}[\gamma]$ be as in Claim 5. By Claim 6, Observation 3.1(ii) holds and so we may assume that there is a cherry in ${\mathcal{F}}^{\prime}$, say $(x,z)$, with $z\in{\color[rgb]{0,0,0}X\setminus\{y\}}$. Now, all of the edges in ${\mathcal{N}}[\gamma]$ must be contained in ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$, otherwise we can apply Lemma 5.5(b) to any edge on ${\mathcal{N}}[\gamma]$ not contained in ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$ and a (C2)(a)(i)-reduction to $x_{i}$ plus induction to see that (CP) holds. But this implies that the edge $e$ in ${\mathcal{B}}$ that contains the vertex adjacent to $x$ and that is not contained in ${\mathcal{N}}[\gamma]$ must be contained in the image of the path in ${\mathcal{F}}^{\prime}$ between $x$ and $z$. Indeed, if this were not the case then, as $e$ must be contained in ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$ by Claim 4, this would imply that $(x,z)$ is not a cherry in ${\mathcal{F}}^{\prime}$. Thus, as $\{x,y\}$ is a component in ${\mathcal{F}}$ and $e$ is not a cut-edge for ${\mathcal{N}}$, we can apply Lemma 5.5(b) to $e$ and the roles of ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ reversed to see that (CP) holds by applying a (C2)(a)(i)-reduction to $x$ and induction.
@@
Claim 8: We can and will assume that there exists some tree ${\mathcal{T}}$ in ${\mathcal{F}}$ that contains a cherry $(x,y)$ with $x,y\in L({\mathcal{B}})$, and such that $|L({\mathcal{B}})\cap L({\mathcal{T}})|\geq 3$.
@@
Now, if ${\mathcal{N}}={\mathcal{B}}$, then since $|L({\mathcal{T}})|\geq 3$ and $|L({\mathcal{B}})|\geq 3$, it immediately follows that $|L({\mathcal{B}})\cap L({\mathcal{T}})|\geq 3$. So, suppose ${\mathcal{N}}\neq{\mathcal{B}}$. Then as $|L({\mathcal{B}})|\geq 3$ it follows that there is some $s\in{\color[rgb]{0,0,0}L({\mathcal{B}})\setminus\{x,y\}}$. If $s\in L({\mathcal{T}})$, then $|L({\mathcal{B}})\cap L({\mathcal{T}})|\geq 3$ and Claim 8 follows. Otherwise $s\not\in L({\mathcal{T}})$ and so there must be some tree ${\mathcal{T}}^{\prime}$ in ${\mathcal{F}}$ with $s\in L({\mathcal{T}}^{\prime})$. Now, since $|L({\mathcal{T}})|\geq 3$, it follows that ${\mathcal{N}}[{\mathcal{T}}]$ must contain the non-pendant cut-edge incident to ${\mathcal{B}}$, since otherwise $|L({\mathcal{B}})\cap L({\mathcal{T}})|\geq 3$. Thus, as ${\mathcal{N}}[{\mathcal{T}}^{\prime}]$ is disjoint from ${\mathcal{N}}[{\mathcal{T}}]$ by the definition of the image of a forest, we must have $L({\mathcal{T}}^{\prime})\subseteq L({\mathcal{B}})$. But ${\mathcal{T}}^{\prime}$ is not a single leaf by Claim 3, and it is not a cherry in the form of an isolated edge by Claim 7. Hence $|L({\mathcal{T}}^{\prime})|\geq 3$, and so ${\mathcal{T}}^{\prime}$ contains a cherry $(r,t)$ with $r,t\in L({\mathcal{B}})$ and $|L({\mathcal{T}}^{\prime})\cap L({\mathcal{B}})|\geq 3$. ∎
@@
Now, by Claim 8, we shall assume that we have a tree ${\mathcal{T}}$ in ${\mathcal{F}}$ that has a cherry $(x,y)$ with $x,y\in L({\mathcal{B}})$ and such that $|L({\mathcal{B}})\cap L({\mathcal{T}})|\geq 3$. Moreover, by Claim 6, Observation 3.1(ii) holds, and so, as well as the cherry $(x,y)$ in ${\mathcal{T}}$, there is a cherry in ${\mathcal{F}}^{\prime}$, say $(x,z)$, with $z\in{\color[rgb]{0,0,0}X\setminus\{y\}}$. Let ${\mathcal{T}}^{\prime}$ denote the tree in ${\mathcal{F}}^{\prime}$ that contains the cherry $(x,z)$. For the cherry $(x,y)$ in ${\mathcal{T}}$ and $\gamma$ the path in ${\mathcal{T}}$ between $x$ and $y$, let $u,v$ and ${\mathcal{N}}[\gamma]$ be as in Claim 5. In addition, let $\delta$ be the path in ${\mathcal{F}}^{\prime}$ between $x$ and $z$ and let ${\mathcal{N}}[\delta]$ be its image in ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$, noting that ${\mathcal{N}}[\delta]$ must have length at least three by Claim 5. We picture this configuration in Figure 5.
@@
Figure 5. A schematic representation illustrating the images ${\mathcal{N}}[\gamma]$ and ${\mathcal{N}}[\delta]$ of the cherries $(x,y)$ and $(x,z)$ in the blob ${\mathcal{B}}$ as well as the associated edges $e$ and $e^{\prime}$ used in the proof of Theorem 6.1. Note that we show that $z\in L({\mathcal{B}})$ in Claim 11.
Claim 9: We can and will assume that every edge in ${\mathcal{N}}[\gamma]$ or in ${\mathcal{N}}[\delta]$ that is in $E({\mathcal{B}})$ must be contained in ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$ or ${\mathcal{N}}[{\mathcal{F}}]$, respectively. In particular, since $x,y\in L({\mathcal{B}})$, ${\mathcal{N}}[\gamma]$ is contained in ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$.
Proof of Claim 9: We consider ${\mathcal{N}}[\gamma]$; the proof for ${\mathcal{N}}[\delta]$ is similar. Suppose $g$ is an edge that is in ${\mathcal{N}}[\gamma]$ and $E({\mathcal{B}})$, but $g$ is not in ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$. Then we can apply a (C2)(a)(i)-reduction to $x$ or $y$, remove the edge $g$ and apply Lemma 5.5(b) to the removed edge plus induction to see that (CP) must hold. Thus, ${\mathcal{N}}[\gamma]$ must be contained in ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$, by Claim 3. ∎
Claim 10: Let $e$ and $e^{\prime}$ be the edges in ${\mathcal{N}}$ that are not in the path ${\mathcal{N}}[\gamma]$ and that contain $u$ and $v$, respectively (see Figure 5). Then we can and will assume that $e$ is in ${\mathcal{N}}[{\mathcal{T}}]$, $e^{\prime}$ is not in ${\mathcal{N}}[{\mathcal{T}}]$, and both $e$ and $e^{\prime}$ are in ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$.
Now, since $e^{\prime}$ is not an edge in ${\mathcal{N}}[{\mathcal{T}}]$, by Claim  4 it follows that $e^{\prime}$ is in ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$. Moreover $e$ is also in ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$, since otherwise as $(x,z)$ is a cherry in ${\mathcal{T}}^{\prime}$ it follows that ${\mathcal{F}}^{\prime}$ contains the component $((x,y),z)$ and so we could apply a (C2)(a)(i)-reduction to $z$, remove $e$ and apply Lemma 5.5(b) to $e$ plus induction to see that (CP) must hold. ∎
Claim 11: We can and will assume that $z\in L(\mathcal{B})$, that every edge in ${\mathcal{N}}[\delta]$ that does not contain $x$ or $z$ is in $E({\mathcal{B}})$, and that ${\mathcal{N}}[\delta]$ is contained in ${\mathcal{N}}[{\mathcal{T}}]$.
@@
Consider the edge $e^{\prime}$ (see Figure 5). Then since $e^{\prime}$ is in ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$ by Claim 10 and it is not a cut-edge, there must be a cherry in ${\mathcal{T}}^{\prime}$, $(r,s)$ say, with $r,s\in L({\mathcal{B}})\setminus\{x,y\}$ (since otherwise ${\mathcal{T}}^{\prime}$ would only have four leaves and we could apply a (C2)(a)(i)-reduction and Lemma 5.5(b) to $e^{\prime}$). But then by Claim 6, without loss of generality ${\mathcal{T}}$ must contain a cherry $(r,a)$, with $a\in X\setminus\{x,y,s\}$. Moreover , since $w$ is a vertex in ${\mathcal{N}}[{\mathcal{T}}]$ we must have $a\in L({\mathcal{B}})$ as every edge in $E({\mathcal{B}})$ and ${\mathcal{N}}[\delta]$ is contained in ${\mathcal{N}}[{\mathcal{T}}]$ by Claim 9. Hence, instead of considering the cherries $(x,y)$ and $(x,z)$ in ${\mathcal{T}}$ and ${\mathcal{T}}^{\prime}$, we can instead consider the cherries $(r,s)$ and $(r,a)$ in ${\mathcal{T}}^{\prime}$ and ${\mathcal{T}}$, respectively, and, reversing the roles of ${\mathcal{T}}$ and ${\mathcal{T}}^{\prime}$, apply our arguments to these cherries instead. In particular, it follows that we may assume $z\in L(\mathcal{B})$.
The last statements in the claim now follow immediately by Claims 3 and 9. ∎
Now, since ${\mathcal{T}}$ contains the cherry $(x,y)$ and $|L({\mathcal{B}})\cap L({\mathcal{T}})|\geq 3$, we can and will assume that either $((x,y),p)$ or $((x,y),(p,q))$ is a pendant subtree of ${\mathcal{T}}$ with $x,y,p,q\in L({\mathcal{B}})$ distinct. Note that by Claim 5, the length of the path ${\mathcal{N}}[\delta]$ is at least three.
@@
Figure 6. The configuration in ${\mathcal{N}}$ that underpins the proofs of Claims 12–14 of Theorem 6.1.
Now, suppose that the length of the path ${\mathcal{N}}[\delta]$ is greater than three. Let $w$, $w^{\prime}$ and $w^{\prime\prime}$ be the vertices in ${\mathcal{N}}[\delta]$ that are distance two, three and four from $x$, respectively. Let $f$, $f^{\prime}$ and $f^{\prime\prime}$ be the edges in ${\mathcal{N}}$ that are not in ${\mathcal{N}}[\delta]$ and that contain the vertices $w$, $w^{\prime}$ and $w^{\prime\prime}$, respectively (see Figure 6).
Claim 12: We can and will assume that $f$ is in ${\mathcal{N}}[{\mathcal{T}}]$.
@@
Proof of Claim 13: Suppose that the claim is not true. Then since $e$ is an edge in ${\mathcal{N}}[{\mathcal{T}}]$ and $f$ is an edge in ${\mathcal{N}}[{\mathcal{T}}]$ by Claim 12, $\{w,w^{\prime}\}$ is in the image of the pendant edge in ${\mathcal{T}}$ that contains $p$, and $((x,y),p)$ is a proper pendant subtree of ${\mathcal{T}}$. Now since every edge in ${\mathcal{N}}[\delta]$ is contained in ${\mathcal{N}}[{\mathcal{T}}]$ by Claim 11, the edge $e^{\prime\prime}$ must be in the image of the pendant edge in ${\mathcal{T}}$ that contains $p$. Moreover, $e^{\prime\prime}$ is not a cut-edge of ${\mathcal{N}}$ since $p\in L({\mathcal{B}})$. so we can apply a (C2)(a)(iii)-reduction to $p$, remove $e^{\prime\prime}$ and apply Lemma 5.5(b) plus induction to see that (CP) holds.∎
@@
Proof of Claim 14: Suppose that the claim is not true. Then using similar arguments to those used in the proof of Claim 13, it follows that without loss of generality, $z=p$, and either $w^{\prime\prime}=z$ and $f^{\prime}$ is in the image of the pendant edge in ${\mathcal{F}}^{\prime}$ that contains $q$, or $f^{\prime}$ is a cut-edge that is not a pendant edge, and $f^{\prime\prime}$ is in the image of the pendant edge in ${\mathcal{F}}^{\prime}$ that contains $q$. In either case, $((x,y),(p,q))$ is a proper pendant subtree of ${\mathcal{T}}$. Furthermore, in either case by removing $f^{\prime}$ or $f^{\prime\prime}$ respectively and applying a (C2)(a)(iii)-reduction to $q$ we can use Lemma 5.5(b) and induction to see that (CP) holds. ∎
@@
7. Discussion
In this paper, we have proven that the hybrid number of two binary phylogenetic trees can be given in terms of cherry picking sequences. There are several interesting future directions of research. For example, in [10] cherry picking sequences for an arbitrary set of rooted, not necessarily binary, phylogenetic trees all having the same leaf set are considered. Is it possible to extend our results to give cherry picking sequences to arbitrary collections of not-necessarily binary phylogenetic trees (or forests) all having the same leaf set? Also, in [10] it was shown that minimum weight cherry picking sequences for collections of rooted phylogenetic trees lead to phylogenetic networks that have a special structure (time consistent and tree-child). Do we get phylogenetic networks with some special structure when we consider cherry picking sequences for unrooted phylogenetic trees?
As mentioned in the introduction, the hybrid number of two unrooted binary phylogenetic trees is equal to the TBR distance between the two trees. Recently, there has been some interest in improving algorithms to compute the TBR distance (see e.g. [18, 20]), and some of the data reductions rules introduced in these papers are similar to the rules that we used to define cherry picking sequences (for example, the reduction described in (C2)(a)(ii) is similar to the (*,3,*)-reduction in [18, Section 3.1]). It would be interesting to see if there is a deeper connection between these two concepts, and to investigate whether, for example, the results presented here could be used to improve depth-bounded search tree algorithms to compute the TBR distance. To investigate this connection, a first step might be to understand whether or not all of the reductions in the definition of a cherry picking sequence are needed, or whether a somewhat simpler list of reductions might be found to define a cherry picking sequence so that Theorem 4.1 still holds.

### v5

@@ 6 Bounding the weight of a cherry picking sequence from above by the hybrid number
Claim 7: Neither ${\mathcal{F}}$ nor ${\mathcal{F}}^{\prime}$ contains a cherry in the form of an isolated edge $\{x,y\}$ with $x,y\in L({\mathcal{B}})$.
@@
Proof of Claim 7: Suppose that ${\mathcal{T}}$ is a cherry in ${\mathcal{F}}$ in the form of the isolated edge $\{x,y\}$ with $x,y\in L({\mathcal{B}})$. Let $u,v,\gamma$ and ${\mathcal{N}}[\gamma]$ be as in Claim 5. By Claim 6, Observation 3.1(ii) holds and so we may assume that there is a cherry in ${\mathcal{F}}^{\prime}$, say $(x,z)$, with $z\in{\color[rgb]{0,0,0}X\setminus\{y\}}$. Now, all of the edges in ${\mathcal{N}}[\gamma]$ must be contained in ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$, otherwise we can apply Lemma 5.6(b) to any edge on ${\mathcal{N}}[\gamma]$ not contained in ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$ and a (C2)(a)(i)-reduction to $x_{i}$ plus induction to see that (CP) holds. But this implies that the edge $e$ in ${\mathcal{B}}$ that contains the vertex adjacent to $x$ and that is not contained in ${\mathcal{N}}[\gamma]$ must be contained in the image of the path in ${\mathcal{F}}^{\prime}$ between $x$ and $z$. Indeed, if this were not the case then, as $e$ must be contained in ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$ by Claim 4, this would imply that $(x,z)$ is not a cherry in ${\mathcal{F}}^{\prime}$. Thus, as $\{x,y\}$ is a component in ${\mathcal{F}}$ and $e$ is not a cut-edge for ${\mathcal{N}}$, we can apply Lemma 5.6(b) to $e$ and the roles of ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ reversed to see that (CP) holds by applying a (C2)(a)(i)-reduction to $x$ and induction. $\blacksquare$
@@
Claim 8: There exists some tree ${\mathcal{T}}$ in ${\mathcal{F}}$ that contains a cherry $(x,y)$ with $x,y\in L({\mathcal{B}})$, and such that $|L({\mathcal{B}})\cap L({\mathcal{T}})|\geq 3$.
@@
Now, if ${\mathcal{N}}={\mathcal{B}}$, then since $|L({\mathcal{T}})|\geq 3$ and $|L({\mathcal{B}})|\geq 3$, it immediately follows that $|L({\mathcal{B}})\cap L({\mathcal{T}})|\geq 3$. So, suppose ${\mathcal{N}}\neq{\mathcal{B}}$. Then as $|L({\mathcal{B}})|\geq 3$ it follows that there is some $s\in{\color[rgb]{0,0,0}L({\mathcal{B}})\setminus\{x,y\}}$. If $s\in L({\mathcal{T}})$, then $|L({\mathcal{B}})\cap L({\mathcal{T}})|\geq 3$ and Claim 8 follows. Otherwise $s\not\in L({\mathcal{T}})$ and so there must be some tree ${\mathcal{T}}^{\prime}$ in ${\mathcal{F}}$ with $s\in L({\mathcal{T}}^{\prime})$. Now, since $|L({\mathcal{T}})|\geq 3$, it follows that ${\mathcal{N}}[{\mathcal{T}}]$ must contain the non-pendant cut-edge incident to ${\mathcal{B}}$, since otherwise $|L({\mathcal{B}})\cap L({\mathcal{T}})|\geq 3$. Thus, as ${\mathcal{N}}[{\mathcal{T}}^{\prime}]$ is disjoint from ${\mathcal{N}}[{\mathcal{T}}]$ by the definition of the image of a forest, we must have $L({\mathcal{T}}^{\prime})\subseteq L({\mathcal{B}})$. But ${\mathcal{T}}^{\prime}$ is not a single leaf by Claim 3, and it is not a cherry in the form of an isolated edge by Claim 7. Hence $|L({\mathcal{T}}^{\prime})|\geq 3$, and so ${\mathcal{T}}^{\prime}$ contains a cherry $(r,t)$ with $r,t\in L({\mathcal{B}})$ and $|L({\mathcal{T}}^{\prime})\cap L({\mathcal{B}})|\geq 3$. $\blacksquare$
@@
Now, by Claim 8, we shall assume that we have a tree ${\mathcal{T}}$ in ${\mathcal{F}}$ that has a cherry $(x,y)$ with $x,y\in L({\mathcal{B}})$ and such that $|L({\mathcal{B}})\cap L({\mathcal{T}})|\geq 3$. Moreover, by Claim 6, Observation 3.1(ii) holds, and so, as well as the cherry $(x,y)$ in ${\mathcal{T}}$, there is a cherry in ${\mathcal{F}}^{\prime}$, say $(x,z)$, with $z\in{\color[rgb]{0,0,0}X\setminus\{y\}}$. Let ${\mathcal{T}}^{\prime}$ denote the tree in ${\mathcal{F}}^{\prime}$ that contains the cherry $(x,z)$. For the cherry $(x,y)$ in ${\mathcal{T}}$ and $\gamma$ the path in ${\mathcal{T}}$ between $x$ and $y$, let $u,v$ and ${\mathcal{N}}[\gamma]$ be as in Claim 5. In addition, let $\delta$ be the path in ${\mathcal{F}}^{\prime}$ between $x$ and $z$ and let ${\mathcal{N}}[\delta]$ be its image in ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$, noting that ${\mathcal{N}}[\delta]$ must have length at least three by Claim 5. We picture this configuration in Fig. 5.
Claim 9: Every edge in ${\mathcal{N}}[\gamma]$ or in ${\mathcal{N}}[\delta]$ that is in $E({\mathcal{B}})$ must be contained in ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$ or ${\mathcal{N}}[{\mathcal{F}}]$, respectively. In particular, since $x,y\in L({\mathcal{B}})$, ${\mathcal{N}}[\gamma]$ is contained in ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$.
Proof of Claim 9: We consider ${\mathcal{N}}[\gamma]$; the proof for ${\mathcal{N}}[\delta]$ is similar. Suppose $g$ is an edge that is in ${\mathcal{N}}[\gamma]$ and $E({\mathcal{B}})$, but $g$ is not in ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$. Then we can apply a (C2)(a)(i)-reduction to $x$ or $y$, remove the edge $g$ and apply Lemma 5.6(b) to the removed edge plus induction to see that (CP) must hold. Thus, ${\mathcal{N}}[\gamma]$ must be contained in ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$, by Claim 3. $\blacksquare$
Claim 10: Let $e$ and $e^{\prime}$ be the edges in ${\mathcal{N}}$ that are not in the path ${\mathcal{N}}[\gamma]$ and that contain $u$ and $v$, respectively (see Fig. 5). Then $e$ is in ${\mathcal{N}}[{\mathcal{T}}]$, $e^{\prime}$ is not in ${\mathcal{N}}[{\mathcal{T}}]$, and both $e$ and $e^{\prime}$ are in ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$.
Now, since $e^{\prime}$ is not an edge in ${\mathcal{N}}[{\mathcal{T}}]$, by Claim  4 it follows that $e^{\prime}$ is in ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$. Moreover $e$ is also in ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$, since otherwise as $(x,z)$ is a cherry in ${\mathcal{T}}^{\prime}$ it follows that ${\mathcal{F}}^{\prime}$ contains the component $((x,y),z)$ and so we could apply a (C2)(a)(i)-reduction to $z$, remove $e$ and apply Lemma 5.6(b) to $e$ plus induction to see that (CP) must hold. $\blacksquare$
@@
Figure 5: A schematic representation illustrating the images ${\mathcal{N}}[\gamma]$ and ${\mathcal{N}}[\delta]$ of the cherries $(x,y)$ and $(x,z)$ in the blob ${\mathcal{B}}$ as well as the associated edges $e$ and $e^{\prime}$ used in the proof of Theorem 6.1. Note that we show that $z\in L({\mathcal{B}})$ in Claim 11.
Claim 11: $z\in L(\mathcal{B})$, every edge in ${\mathcal{N}}[\delta]$ that does not contain $x$ or $z$ is in $E({\mathcal{B}})$, and ${\mathcal{N}}[\delta]$ is contained in ${\mathcal{N}}[{\mathcal{T}}]$.
@@
Consider the edge $e^{\prime}$ (see Fig. 5). Then since $e^{\prime}$ is in ${\mathcal{N}}[{\mathcal{F}}^{\prime}]$ by Claim 10 and it is not a cut-edge, there must be a cherry in ${\mathcal{T}}^{\prime}$, $(r,s)$ say, with $r,s\in L({\mathcal{B}})\setminus\{x,y\}$ (since otherwise ${\mathcal{T}}^{\prime}$ would only have four leaves and we could apply a (C2)(a)(i)-reduction and Lemma 5.6(b) to $e^{\prime}$). But then by Claim 6, without loss of generality ${\mathcal{T}}$ must contain a cherry $(r,a)$, with $a\in X\setminus\{x,y,s\}$. Moreover , since $w$ is a vertex in ${\mathcal{N}}[{\mathcal{T}}]$ we must have $a\in L({\mathcal{B}})$ as every edge in $E({\mathcal{B}})$ and ${\mathcal{N}}[\delta]$ is contained in ${\mathcal{N}}[{\mathcal{T}}]$ by Claim 9. Hence, instead of considering the cherries $(x,y)$ and $(x,z)$ in ${\mathcal{T}}$ and ${\mathcal{T}}^{\prime}$, we can instead consider the cherries $(r,s)$ and $(r,a)$ in ${\mathcal{T}}^{\prime}$ and ${\mathcal{T}}$, respectively, and, reversing the roles of ${\mathcal{T}}$ and ${\mathcal{T}}^{\prime}$, apply our arguments to these cherries instead. In particular, it follows that we may assume $z\in L(\mathcal{B})$.
The last statements in the claim now follow immediately by Claims 3 and 9. $\blacksquare$
Now, since ${\mathcal{T}}$ contains the cherry $(x,y)$ and $|L({\mathcal{B}})\cap L({\mathcal{T}})|\geq 3$, we will assume that either $((x,y),p)$ or $((x,y),(p,q))$ is a pendant subtree of ${\mathcal{T}}$ with $x,y,p,q\in L({\mathcal{B}})$ distinct. Note that by Claim 5, the length of the path ${\mathcal{N}}[\delta]$ is at least three.
@@
Figure 6: The configuration in ${\mathcal{N}}$ that underpins the proofs of Claims 12–14 of Theorem 6.1.
Now, suppose that the length of the path ${\mathcal{N}}[\delta]$ is greater than three. Let $w$, $w^{\prime}$ and $w^{\prime\prime}$ be the vertices in ${\mathcal{N}}[\delta]$ that are distance two, three and four from $x$, respectively. Let $f$, $f^{\prime}$ and $f^{\prime\prime}$ be the edges in ${\mathcal{N}}$ that are not in ${\mathcal{N}}[\delta]$ and that contain the vertices $w$, $w^{\prime}$ and $w^{\prime\prime}$, respectively (see Fig. 6).
Claim 12: $f$ is in ${\mathcal{N}}[{\mathcal{T}}]$.
@@
Proof of Claim 13: Suppose that the claim is not true. Then since $e$ is an edge in ${\mathcal{N}}[{\mathcal{T}}]$ and $f$ is an edge in ${\mathcal{N}}[{\mathcal{T}}]$ by Claim 12, $\{w,w^{\prime}\}$ is in the image of the pendant edge in ${\mathcal{T}}$ that contains $p$, and $((x,y),p)$ is a proper pendant subtree of ${\mathcal{T}}$. Now since every edge in ${\mathcal{N}}[\delta]$ is contained in ${\mathcal{N}}[{\mathcal{T}}]$ by Claim 11, the edge $e^{\prime\prime}$ must be in the image of the pendant edge in ${\mathcal{T}}$ that contains $p$. Moreover, $e^{\prime\prime}$ is not a cut-edge of ${\mathcal{N}}$ since $p\in L({\mathcal{B}})$. so we can apply a (C2)(a)(iii)-reduction to $p$, remove $e^{\prime\prime}$ and apply Lemma 5.6(b) plus induction to see that (CP) holds. $\blacksquare$
@@
Proof of Claim 14: Suppose that the claim is not true. Then using similar arguments to those used in the proof of Claim 13, it follows that without loss of generality, $z=p$, and either $w^{\prime\prime}=z$ and $f^{\prime}$ is in the image of the pendant edge in ${\mathcal{F}}^{\prime}$ that contains $q$, or $f^{\prime}$ is a cut-edge that is not a pendant edge, and $f^{\prime\prime}$ is in the image of the pendant edge in ${\mathcal{F}}^{\prime}$ that contains $q$. In either case, $((x,y),(p,q))$ is a proper pendant subtree of ${\mathcal{T}}$. Furthermore, in either case by removing $f^{\prime}$ or $f^{\prime\prime}$ respectively and applying a (C2)(a)(iii)-reduction to $q$ we can use Lemma 5.6(b) and induction to see that (CP) holds. $\blacksquare$
@@
7 Discussion
In this paper, we have proven that the hybrid number of two binary phylogenetic trees can be given in terms of cherry picking sequences. There are several interesting future directions of research. For example, Humphries et al. [11] have considered a certain type of cherry picking sequence to characterize arbitrary size collections of rooted phylogenetic trees that can be displayed by a rooted phylogenetic network that is time consistent and tree-child. It would be interesting to investigate extensions of our results to computing the hybrid number of an arbitrary collection of not necessarily binary phylogenetic trees (or forests) that can be displayed by an unrooted phylogenetic network of a given class such as tree-child or orchard (see [8] for the definition of such networks).
As mentioned in the introduction, the hybrid number of two unrooted binary phylogenetic trees is equal to the TBR distance between the two trees. Recently, there has been some interest in improving algorithms to compute the TBR distance (see e.g. [19, 20]), and some of the data reductions rules introduced in these papers are similar to the rules that we used to define cherry picking sequences (for example, the reduction described in (C2)(a)(ii) is similar to the (*,3,*)-reduction in [19, Section 3.1]). It would be interesting to see if there is a deeper connection between these two concepts, and to investigate whether, for example, the results presented here could be used to improve depth-bounded search tree algorithms to compute the TBR distance. To investigate this connection, a first step might be to understand whether or not all of the reductions in the definition of a cherry picking sequence are needed, or whether a somewhat simpler list of reductions might be found to define a cherry picking sequence so that Theorem 4.1 still holds.